A container of mass 200kg rests on the back of an open truck.if the truck accelerates 1.5m/secondsquare.what is the minimum co-efficient of static friction between the container and the bed of the truck required to prevent the container from sliding off the back of the truck is?

Grade:10

1 Answers

Aman Bansal
11 years ago

Dear Likhhit,

Fsf = ma, Fsf = k * m * g.
k * m * g = m * a
k * g = a
k = a / g
k = 1.5 / 9.8 = 0.1531

Cracking IIT just got more exciting,It s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple  to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert